OpenStudy (anonymous):

The equations of two lines are x - 3y = 6 and y = 3x + 2. Determine if the lines are parallel, perpendicular or neither.

OpenStudy (anonymous):

is it perpendicular

OpenStudy (anonymous):

The lines are perpendicular to each other. When you solve x-3y=6 you will get -3y=x+6 .... when you divide both sides by -3 ... you will get Y=-X/3 - 2. The slope is flipped from the first. The slope of line 1 is Y=3/1x The second line is -X/3 ..... the slope is flipped. When you flip the slope ... the slope becomes a perpendicular line. Therefore, these lines are PERPENDICULAR to each other.

OpenStudy (***[isuru]***):

nope>>>> @rbeckford134 u have missed something... they r not perpendicular to each other x - 3y = 6 3y = x -6 y = x/3 -2 so @rbeckford134 u r wrong :-(

OpenStudy (anonymous):

@ISURU .... how are you getting positive 3? Its a -3y ..... The value X transfers across the = sign to become a -x not a positive. Remember that X stands place of 1 ... there for ... -X transfers over to become a -1.

OpenStudy (anonymous):

Use the equation of a line .... Y = MX + B. @ISURU

OpenStudy (anonymous):

When you graph the lines .... this is what you get:

OpenStudy (***[isuru]***):

hey ... watch the function of the line given in the question... it's x - 3y = 6